KEK

HIGH ENERGY ACCELERATOR RESEARCH ORGANIZATION 1-1 OHO TSUKUBA-SHI IBARAKI-KEN 305-0801, JAPAN

TELEPHONE: 0298-64-1171 FAX: 0298-64-2580

November 30, 2000

Prof. William Barletta, Lawrence Berkeley Laboratory, MS 50-149, 1 Cyclotron Road, Berkeley, CA 94720, U.S.A.

Dear Prof. Barletta,

Enclosed you will find the manuscript of our paper entitled

”Lorentz Angle Measurement for CO_{2}/Isobutane Gas Mixtures”

submitted to Nuclear Instruments and Methods in Physics Research Section A.

You will also find a set of original figures used in the paper and three copies of the manuscript. We are grateful if you kindly consider our paper for publication in your journal.

My mail address is

Keisuke Fujii IPNS, KEK 1-1, Oho, Tsukuba Ibaraki 305-0801, Japan E-mail address: fujiik@jlcuxf.kek.jp

to which further correspondence should be made.

Thank you very much in advance.

Sincerely yours,

Keisuke Fujii